Urmamaisnudad292:

Help me please :) 1. 2x^4+5x^3+x-1 by x^2-2x+1 2. 2x^3+x^2+4 by x-3 Solve using polynomial long division and show work

wio:

Do you know how to do polynomial long division?

Urmamaisnudad292:

No

wio:

Would you like to learn how to do it?

Urmamaisnudad292:

If you wanna teach me

wio:

Alright, let's start with \((2x^4+5x^3+x-1) / (x^2-2x+1) \).

wio:

We start with the first terms. \(2x^4/x^2 = ?\)

wio:

Could you do this division?

wio:

So \(2x^4/x^2 = 2x^2\).

Urmamaisnudad292:

Okay

wio:

Can you multiply polynomials? Now we multiply \(2x^2\times (x^2-2x+1)\).

wio:

\[ 2x^2\times(x^2-2x+1) = 2x^4-4x^3+2x^2 \]

wio:

We subtract this from the thing being divided. \[ 2x^4+5x^3+x-1 - (2x^4-4x^3+2x^2) = 9x^3-2x^2+x-1 \]

wio:



wio:

Basically what I've done so far is the first step.

Urmamaisnudad292:

Ooo okay

wio:

Since the divisor starts with \(x^2\), and the remaining part starts with \(9x^3\), we can still keep going. Now we do \(9x^3/x^2 = 9x\).

wio:

\(9x(x^2-2x+1) = 9x^3 -18x^2+9x\)

wio:

\[ 9x^3-2x^2+x-1 - ( 9x^3 -18x^2+9x) = 16x^2-8x-1 \]

wio:

It keeps going like this. Does this make sense?

wio:



wio:

We can still keep going since the starting parts are \(x^2\) and \(16x^2\).

wio:

\(16x^2/x^2=16\). \(16(x^2-2x+1) = 16x^2-32x+16\). \(16x^2-8x+1 - (16x^2-32x+16) = 24x-15 \).

wio:

At this point we have to stop, because the starting parts are \(x^2\) and \(24x\). Since the divisor has a larger exponent, we simply accept it is a remainder. Thus we end with \[ \frac{24x-15}{x^2-2x+1} \]. If we add this to the parts we already divided out earlier, the final answer is: \[ 2x^2+9x+16+\frac{24x-15}{x^2-2x+1} \]

wio:

The algorithm here is: Step 1: Take the dividend's largest term and divide by the divisors largest term. Step 2: Multiply Step 1's result with the divisor. Step 3: Subtract Step 2's result from the dividend. Continue until the dividend's largest term is lower than the divisors largest term.

wio:

I think it would be great if you tried to do question 2 on your own, but I can help if you still don't get it.

Urmamaisnudad292:

Okay thank you so much <33

wio:

https://www.mathsisfun.com/algebra/polynomials-division-long.html This website also explains it a bit.

Urmamaisnudad292:

Thank you bb